1/x-4 + 5 = 3/x-4. I need to know how to solve this type of equation.
Answer by usyim88hk(157)

First we should multiply both sides by (x-4) so that we won't have any fraction on both sides of the equation:
:
(x-4)(1/x-4) + (x-4)5 = (3/x-4)(x-4)
1 + 5(x-4) = 3
1 + 5x - 20 = 3
5x = 22
x = 22/5
x = 4.4